use the following weighted to recalculate each student's average

English 3

math 3

chemistry 2

history 2

phys.Ed 1


Nick:

Zach

overall who has the higher mark

calculate a 10% trimmed mean for the data set

consider your values for the mean and the trimmed mean. which statistics is a better indicator of how most students will perform on the test? explain you choise


1
Expert's answer
2021-03-31T08:59:14-0400

Average= Total weighted divided by 2

Average=(3+3+2+2+1) divided by 2

Average=11/2=5.5

"0.1 trimmed mean= 0.9\times 5.5=4.95"

Trimmed mean is better indicator because it caters for outliers and errors